Ok wow I need some major help!! I need help on how to solve ordered pairs. I looked in my text book and that made me more confused!!
A) (-9,7), 7x+y<-9 What is the solution of the ordered pair?
B) 2x+9y=21, 5x-9y=-42
C) 2x+y=5, 3x+4y=-15
D) x=8, y=-7
E) Solve by te substitution method: 9m+n=8, m-9n=10

Answer by stanbon(75887)

Ok wow I need some major help!! I need help on how to solve ordered pairs. I looked in my text book and that made me more confused!!
A) (-9,7), 7x+y<-9 What is the solution of the ordered pair?
You are asked to determine whether (-9,7) gives you a true statement.
Substitute to get 7*-9+7<-9; -63+7<-9; -56 < -9
That is true.
-----------------------------------
B) 2x+9y=21,
5x-9y=-42
--
Add the two equations to get:
7x = -21
x = -3
Substitute to solve for "y":
==================================
C) 2x+y=5
3x+4y=-15
---
Multiply thru the 1st equation by 4 to get:
8x+4y = 20
---
Subtract the 2nd equation to get:
5x = 35
x = 7
Substitute to solve for "y".
==================================
D) x=8, y=-7
Hummmmm????
==================================
E) Solve by the substitution method:
9m+n=8,
m-9n=10
---
m = 9n+10
Substitute for m to solve for "n":
9(9n+10)+n = 8
82n + 90 = 8
82n = -82
n = -1
---
Substitute to solve for "m"
================================
Cheers,
Stan H.